d. Full production phase 2. Which is statement i FALSE df lominy test a. The results from the test will indicate if a material can be successfully case hardened b. The metal hardness can be measured by a durometer c. The resulting hardenability curve plots hardness versus distance from the d. A metal sample is cooled quickly by submerging in water Machi inin quenched end 3. A 30mm diameter piece of bar stock in a lathe has a surface speed of 200mms. What a. 2.1 rpm b. 127.3 rpm c. 13.3 rpm d. Not possible to calculate What is its rpm? 4 You are buying a machine that costs $250,000. You will pay for the machine in equal m monthly payments monthly payment? 6years. The interest rate is 12% per year compounded monthly What is the a. $5,561 11 b. $4,166.67 c. $30,033.46 d. None of the above 5. Which of the following materials has isotropic bulk mechanical properties? a. Single crystal diamond b. Fine grained polycrystalline copper c. Single crystal silicon d. A cylindrical ceramic specimen 6. which of the following materials is expected to undergo the least strain in a uniaxial tension test with a load of 1MPa? a. Copper b Steel c. Silicone rubber d. Aluminum 7. Which of the following processes is NOT part of the manufacture of silicon wafer substrates? a. Crystal growth b. Boule slicing c. Polishing d. EDM 8. Which of the following processes can be performed on a lathe: a. Boring b. Pocketing c. Broaching d. None of the above EGR 2180 Page 3 gust 2018
Answer #1

2. b..... the metal hardness can be measured by durometer

3. b......... n=200/(3.14*30) n=2.1 rps=2.1*60=127.3 rpm

4. b..... considering time value money it is 4166.67$

5. b......fine grained polycrystaline copper

6. b......steel

7. b......boule slicing

8. a......boring